### v1

# Adversarially Perturbed Precision Matrix Estimation

Abstract
Precision matrix estimation is a fundamental topic in multivariate statistics and modern machine learning. This paper proposes an adversarially perturbed precision matrix estimation framework, motivated by recent developments in adversarial training. The proposed framework is versatile for the precision matrix problem since, by adapting to different perturbation geometries, the proposed framework can not only recover the existing distributionally robust method but also inspire a novel moment-adaptive approach to precision matrix estimation, proven capable of sparsity recovery and adversarial robustness. Notably, the proposed perturbed precision matrix framework is proven to be asymptotically equivalent to regularized precision matrix estimation, and the asymptotic normality can be established accordingly. The resulting asymptotic distribution highlights the asymptotic bias introduced by perturbation and identifies conditions under which the perturbed estimation can be unbiased in the asymptotic sense. Numerical experiments on both synthetic and real data demonstrate the desirable performance of the proposed adversarially perturbed approach in practice.
1 Introduction
The precision matrix, also known as the inverse covariance matrix, can reveal the conditional correlations between pairs of variables given all other variables (Fan et al., 2016). Estimating the precision matrix is essential in many problems, including neuroimaging (Huang et al., 2010), network analysis (Torri et al., 2018), portfolio optimization (Procacci and Aste, 2022), and genomics (Schäfer and Strimmer, 2005), among others.
Suppose $\bm{x}_{1},\ldots,\bm{x}_{n}$ are $n$ independent and identically distributed samples drawn from a $d$-dimensional multivariate normal distribution. For simplicity, we assume that the population mean is zero and the population covariance matrix $\Sigma$ is positive definite and unknown. The standard approach to obtain the precision matrix estimation, i.e., $C=\Sigma^{-1}$, is to identify the minimizer of the negative log-likelihood function:
$\min_{C\succ 0}-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}^{\top}C\bm{x}],$ (1)
where $C\succ 0$ means that $C$ is positive definite, $\mathrm{det}C$ denotes the determinant of the matrix $C$, $\mathbb{P}_{n}$ is the empirical distribution. When the matrix dimension is small such that $d<n$, the solution to the maximum likelihood estimation problem is the inverse of the sample covariance matrix $\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}\bm{x}^{\top}]$. However, if the matrix dimension is large such that $d>n$, the maximum likelihood estimation problem does not admit a solution. Numerous studies have been conducted to explore shrinkage and regularization frameworks for precision matrix estimation to achieve estimations which are invertible, well-conditioned, and enjoy desirable statistical properties such as sparsity (Yuan and Lin, 2007; Ledoit and Wolf, 2012; Nguyen et al., 2022). We will show later that the proposed adversarially perturbed framework in this paper can recover some of these shrinkage and regularized effects by selecting different types of perturbations.
Our proposed adversarially perturbed precision matrix framework is inspired by recent developments in trustworthy machine-learning—adversarial training (Goodfellow et al., 2015). The perturbed estimation is obtained by solving the following optimization problem:
$\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\max_{\|\bm{\Delta}\|\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})\right]\right\},$ (2)
where $\|\cdot\|$ is the norm of the perturbation $\bm{\Delta}$, and $\delta$ is the magnitude of the perturbation. Instead of minimizing the empirical average of the likelihood function, the adversarially perturbed estimation problem (2) seeks for the minimizer of the loss under the most adversarial perturbation $\bm{\Delta}$, which is determined by both $C$ and $\bm{x}$ through the inner maximization problem in (2), subject to the constraint that the norm of the perturbation is bounded.
In this paper, we will focus on the $\ell_{p}$-norm of the perturbation. When $p=2$, it can be proven that the perturbed estimation problem (2) is equivalent to the Wasserstein shrinkage estimation proposed in Nguyen et al. (2022). The resulting optimization problem is tractable and yields a nonlinear shrinkage estimator that is well-conditioned, rotation equivalent, and preserves the eigenvalue order (Nguyen et al., 2022). When $p=\infty$, by optimizing a convex upper-bound surrogate of the objective function in the outer minimization in problem (2), the resulting estimation can be proven to promote sparsity and adversarial robustness.
We demonstrate that introducing perturbations is asymptotically equivalent to imposing regularization in precision matrix estimation, provided the perturbation magnitude is small. The corresponding first-order and second-order regularization terms are also given. Based on the regularization effect, the asymptotic distribution of the adversarially perturbed estimation is derived. The asymptotic distribution characterizes the bias induced by the inner maximization problem in our proposed framework (2). We suppose the perturbation magnitude follows the rule $\delta=\eta/n^{\gamma}$, where $\eta>0$ and $\gamma>0$. Our analysis yields the following two regimes: When $\gamma>1/2$, the asymptotic bias is zero. In this regime, the adversarial perturbation decays sufficiently fast such that it does not affect estimation accuracy. When $\gamma\leq 1/2$, the asymptotic distribution has a non-zero bias. Furthermore, specifically for $\ell_{\infty}$-norm perturbations using the upper-bound surrogate of the objective function, the derived asymptotic distribution demonstrates that sparsity recovery is achievable when $\gamma=1/2$. The regularization reformulation and asymptotic distribution provide insights into the statistical behavior of both Wasserstein shrinkage estimation and the sparse estimation recovered by our novel framework. See a summary table of the associated asymptotic behavior in Table 1, Section 5.
Numerical experiments are delivered for the case when $p=\infty$ and the surrogate objective is considered. We carry out the numerical experiments on both synthetic data and a real-world dataset. Firstly, the numerical experiments on the synthetic data validate our theoretical results that the perturbed estimation can recover the sparsity and achieve adversarial robustness. Secondly, by adopting linear discriminant analysis based on precision matrix estimation for the real dataset, our method shows more desirable practical performance.
1.1 Related Work
1.1.1 Adversarial Training
As the modern artificial intelligence and machine learning models grow more powerful and widely used, their reliability becomes increasingly important. One major challenge to reliability is that machine learning models can be vulnerable to small changes in input data, which may lead to unreliable outputs (Goodfellow et al., 2015). To improve the robustness of the models, adversarial training has been introduced as a principled approach by optimizing against adversarial perturbations posed in the input data (Goodfellow et al., 2015; Madry et al., 2018).
Different from the traditional empirical risk minimization approach, the resulting adversarial training problem is a min-max robust optimization problem and thus introduces a new statistical estimation paradigm. While most of the relevant existing work focus on the vector-valued parameter (Goodfellow et al., 2015; Madry et al., 2018; Javanmard and Soltanolkotabi, 2022; Ribeiro et al., 2023; Xie and Huo, 2024, 2025), our proposed framework provides novel insights on how the adversarial training motivates a new estimation framework for the matrix-valued parameter. The proposed framework serves as an alternative for practitioners when the precision matrix estimation with desirable properties such as sparsity and adversarial robustness is needed.
Adversarial training is closely related to the distributionally robust optimization. It is well known that adversarial training can be viewed as a special case of the Wasserstein distributionally robust optimization (Staib and Jegelka, 2017; Gao et al., 2024). This paper, however, provides another perspective by demonstrating that, in the context of precision matrix estimation, the Wasserstein distributionally robust optimization is a special case of adversarial training.
1.1.2 Precision Matrix Estimation
Precision matrix estimation has been extensively studied under different settings. Classical approaches primarily focus on enforcing sparsity via $\ell_{1}$-regularization or $\ell_{1}$-minimization (Yuan and Lin, 2007; Friedman et al., 2008; Cai et al., 2011). As applications have broadened, researchers have developed methods robust to data anomalies. For instance, Loh and Tan (2018) establishes a framework robust to data contamination, Tang et al. (2021) focuses on mitigating the impact of outliers, Avella-Medina et al. (2018) studies robustness against heavy-tailed distributions, and the method framed in Nguyen et al. (2022) addresses distributional shifts.
Our framework can recover the Wasserstein shrinkage estimation (Nguyen et al., 2022) under $\ell_{2}$-perturbation while being capable of the sparsity recovery achieved by classical $\ell_{1}$-regularized methods (Yuan and Lin, 2007; Friedman et al., 2008). In addition, by defining the model through an adversarial training lens, we achieve intrinsic adversarial robustness. Specifically, as detailed in Section 6.1, under $\ell_{\infty}$-perturbation and a specific surrogate objective, our formulation reduces to a moment-adaptive weighted $\ell_{1}$-regularized precision matrix estimation problem, thereby simultaneously achieving sparse recovery and adversarial robustness.
1.2 Notations and Definitions
$\bm{x}]_{i}$ denotes the $i$th component of vector $\bm{x}$. $\operatorname{sign}([\bm{x}]_{i})$ denotes the sign of $\bm{x}]_{i}$ (where we let $\operatorname{sign}(0)=0$), and accordingly we let $\operatorname{sign}(\bm{x})=(\operatorname{sign}([\bm{x}]_{1}),...,\operatorname{sign}([\bm{x}]_{d}))^{\top}$. We use $\|\bm{x}\|_{p}$ to denote the $\ell_{p}$-norm of the vector $\bm{z}$, i.e., $\|\bm{x}\|_{p}^{p}=\sum_{j=1}^{d}|x_{j}|^{p},1\leq p<\infty,\|\bm{x}\|_{\infty}=\max_{1\leq j\leq d}|x_{j}|.$ $\odot$ denotes the component-wise product. $\mathbb{I}(\cdot)$ denotes the indicator function. $\mathrm{vec}(X)$ denotes the vectorization of the matrix $X$.
1.3 Organization of this Paper
The remainder of this paper is organized as follows. In Section 2, we study the adversarially $\ell_{2}$-perturbed precision matrix estimation and its equivalence to the Wasserstein shrinkage estimation. In Section 3, we investigate the formulation and the properties of the adversarially $\ell_{\infty}$-perturbed precision matrix estimation. In Section 4, we introduce the regularization effect of the proposed framework. In Section 5, we study the asymptotic distribution of the proposed precision matrix estimation. Numerical experiments are conducted and analyzed in Section 6. Future work is discussed in Section 7. The proofs are relegated to the Supplementary Material whenever possible.
2 Adversarially $\ell_{2}$-perturbed Precision Matrix Estimation
In this section, we will discuss the $\ell_{2}$-perturbed precision matrix estimation.
The adversarially perturbed precision matrix estimation under $\ell_{2}$-norm is the solution to the following optimization problem:
$\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\max_{\|\bm{\Delta}\|_{2}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})\right]\right\},$ (3)
where the perturbation norm in the constraint set in the inner maximization problem is chosen as $\ell_{2}$ norm.
2.1 Tractable Reformulation
In this subsection, we give the convex tractable reformulations of the adversarially $\ell_{2}$-perturbed precision matrix estimation.
Theorem 2.1 (Convex Reformulations under $\ell_{2}$-perturbation).
The problem (3) is equivalent to the following problem:
$\inf_{C\succ 0,\lambda I-C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}^{\top}C\bm{x}]+\lambda\delta^{2}+\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}^{\top}C(\lambda I-C)^{-1}C\bm{x}]\right\},$ (4)
which further admits the following reformulation:
$\inf_{\lambda,\{t_{i}\},C}$ $-\log\mathrm{det}C+\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}^{\top}C\bm{x}_{i}+\lambda\delta^{2}+\frac{1}{n}\sum_{i=1}^{n}t_{i}$ (5)
$\mathrm{subject\ to}$ $C\succ 0,\lambda I-C\succ 0,$
$\begin{pmatrix}t_{i}&\bm{x}_{i}^{\top}C\\ C\bm{x}_{i}&\lambda I-C\end{pmatrix}\succeq 0,\quad\forall i=1,\dots,n.$
The reformulation (4) is obtained by deriving the dual of the inner maximization problem in (3), which admits a closed-form solution under $\ell_{2}$-perturbation. Subsequently, applying the Schur complement yields formulation (5).
Problem (5) is a convex semidefinite optimization problem with a log-determinant barrier term in the objective function and linear matrix inequalities in the decision variables $\lambda,\{t_{i}\},C$. Such problems fall within the framework of semidefinite (conic) programming and can be solved efficiently by interior-point methods (Boyd and Vandenberghe, 2004; Nemirovski, 2007).
2.2 Equivalence to the Wasserstein Shrinkage Estimation
In this subsection, we will show that the $\ell_{2}$-perturbed precision matrix estimation problem is equivalent to the Wasserstein distributionally robust precision matrix estimation proposed in Nguyen et al. (2022), which they coin as the Wasserstein shrinkage estimation.
The optimization problem for obtaining the Wasserstein shrinkage estimation introduced in Nguyen et al. (2022) can be written as the following problem:
$\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\sup_{S\in\mathcal{U}_{\bar{A},\rho}}\mathrm{tr}(SC),\right\},$ (6)
where $\mathcal{U}_{\bar{A},\rho}$ is the uncertainty set and $\rho$ is the radius of the set. The uncertainty set is centered at the sample covariance $\bar{A}=\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}\bm{x}^{\top}]$ and is defined in terms of the metric induced by the 2-Wasserstein distance on the family of normal distributions, and we have that
$\mathcal{U}_{\bar{A},\rho}=\left\{S\succeq 0:\mathrm{tr}(\bar{A})+\mathrm{tr}(S)-2\mathrm{tr}(\sqrt{\bar{A}^{\frac{1}{2}}S\bar{A}^{\frac{1}{2}}})\leq\rho^{2}\right\}.$
Nguyen et al. (2022) shows that the min-max optimization problem in (6) admits a convex reformulation, as restated in Theorem 2.2.
Theorem 2.2 (Nguyen et al. (2022)).
For any fixed $\rho>0$, the Wasserstein distributionally robust estimation problem (6) is equivalent to
$\inf_{C\succ 0,\lambda I-C\succ 0}\left\{-\log\mathrm{det}C+\lambda\left(\rho^{2}-\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}\bm{x}_{i}^{\top}\right)+\lambda^{2}\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}(\lambda I-C)^{-1}\bm{x}_{i}^{\top}\right\}.$ (7)
The equivalence of our adversarially perturbed estimation problem (3) and distributionally robust trained problem (6) can be built upon problem (4) and (7) based on algebraic calculations. We state this result in the following theorem.
Theorem 2.3 (Equivalence).
For fixed $\rho=\delta$, the $\ell_{2}$-perturbed precision matrix estimation problem (3) is equivalent to the Wasserstein distributionally robust precision matrix estimation problem (6).
Theorem 2.3 implies that, in the precision matrix estimation problem, our proposed adversarially perturbed approach could recover the Wasserstein distributionally estimator if the $\ell_{2}$-perturbation is chosen. Imposing an $\ell_{2}$-perturbation to the observation $\bm{x}$ is equivalent to constructing an ambiguity set for the sample covariance matrix in terms of the Wasserstein distance. As illustrated in Nguyen et al. (2022), the resulting estimator is invertible, rotation equivalent, preserves the eigenvalue order, and can be obtained by solving a sequential quadratic approximation algorithm.
Although arriving at the same solution, our proposed perturbed framework (3) adopts a sample-wise perturbation perspective. Crucially, this formulation admits a tractable asymptotic expansion, from which the asymptotic distribution is derived, seeing Section 5. This framework elucidates the statistical effect of shrinkage and enables statistical inference.
3 Adversarially $\ell_{\infty}$-perturbed Precision Matrix Estimation
In this section, we explore the adversarially $\ell_{\infty}$-perturbed precision matrix estimation.
For the adversarially perturbed precision matrix estimation (2), if the norm of the perturbation is chosen as $\ell_{\infty}$-norm, we have the following formulation:
$\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\max_{\|\bm{\Delta}\|_{\infty}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})\right]\right\}.$ (8)
3.1 Tractable Surrogate
Notably, the inner maximization problem in problem (8) is NP-hard (Nemirovski et al., 1999). To ensure computational tractability, we replace the NP-hard inner maximization in (8) with a convex upper-bound surrogate. Based on the triangle inequality on the perturbation terms, we derive the following formulation:
$\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}C\bm{x}\right]+2\delta\sum_{j=1}^{d}\sum_{k=1}^{d}\omega_{k}|[C]_{kj}|+\delta^{2}\|C\|_{1,1}\right\}$ (9)
where
$\omega_{k}=\mathbb{E}_{\mathbb{P}_{n}}\left[\left|[\bm{x}]_{k}\right|\right].$
The formulation (9) can be interpreted as a regularized precision matrix estimation with penalties with orders $O(\delta)$ and $O(\delta^{2})$. The term $2\delta\sum_{j,k}\omega_{k}|[C]_{kj}|$ scales with $O(\delta)$ and functions as a moment-adaptive weighted $\ell_{1}$-regularizer. The weights $\omega_{k}$’s enforce a heterogeneous shrinkage scheme where entries associated with variables of larger average magnitude are penalized more heavily. Since high-magnitude features act as high-leverage points for adversarial exploitation, this weighting scheme promotes an estimation that is not only sparse but also intrinsically robust to adversarial manipulation. The quadratic term $\delta^{2}\|C\|_{1,1}$ addresses the second-order interactions of the adversarial perturbation. By penalizing the entry-wise magnitude of matrix $C$, this quadratic regularizer serves as a global stability adjustment, ensuring the surrogate remains a rigorous upper bound for the original objective in (8).
3.1.1 Properties
We will discuss the properties of the estimation obtained from (9), including the adversarial robustness stated in Proposition 3.1 and the sparsity recovery stated in Corollary 3.2.
Proposition 3.1 (Adversarial Robustness).
Suppose $\widetilde{C}$ is the minimizer of problem (9), and $\mathcal{S}(\cdot)$ is the associated objective function. Then, the adversarial robust loss of $\widetilde{C}$ is upper bounded, i.e.,
$\log\mathrm{det}\widetilde{C}+\mathbb{E}_{\mathbb{P}_{n}}\left[\max_{\|\bm{\Delta}\|_{p}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}\widetilde{C}(\bm{x}+\bm{\Delta})\right]\leq\inf_{C\succ 0}\mathcal{S}(C),\quad 1\leq p\leq\infty.$
Proposition 3.1 implies that the adversarially robust loss of the estimation $\widetilde{C}$ obtained from problem (9) under $\ell_{p}$-perturbation, which is defined as the worst-case loss under adversarial perturbations posed to the samples, is upper bounded. This property is due to the nature of adversarial training, the motivation behind our proposed problem, and the way we formulate problem (9) as an upper-bound surrogate.
Corollary 3.2 (Sparsity Recovery).
Let $\widetilde{C}$ be the minimizer of problem (9). Suppose that $\delta=\eta/\sqrt{n}$ with $\eta>0$. Then, as $n\to\infty$, the asymptotic distribution of $\widetilde{C}]_{ij}$ has a positive probability mass at 0 whenever $\Sigma^{-1}]_{ij}=0$.
Remark 3.3.
Corollary 3.2 follows immediately from Theorem 5.2 in Section 5 and Proposition 3.8 in Xie and Huo (2025).
Corollary 3.2 implies that for zero entries in the true precision matrix $\Sigma^{-1}$, the asymptotic distribution of the corresponding estimates from (9) has a positive probability at zero. We call this property sparsity recovery, which is theoretically guaranteed if the adversarial perturbation order is $1/\sqrt{n}$. The sparsity is induced by the first penalty term in (9). Notably, while similar to the shrinkage achieved in $\ell_{1}$-regularized precision matrix estimation (Yuan and Lin, 2007; Friedman et al., 2008), this term incorporates the adaptive adjustment for adversarial robustness discussed earlier
4 Regularization Effect
We discuss the regularization effect induced by the adversarial perturbation in the proposed framework.
We characterize this regularization effect mathematically in the following proposition.
Proposition 4.1 (Regularization Effect).
Suppose the random variable $\bm{x}$ follows a distribution $\bm{P}$ with $\mathbb{E}_{\bm{P}}[\|\bm{x}\|_{p}]<\infty$ and $\bm{P}([C\bm{x}]_{i}=0)=0$, $C$ is a positive definite matrix with $\|C\|_{p\to q}<\infty$, and $p\in(1,\infty]$ with $1/p+1/q=1$. Then, we have that
$\mathbb{E}_{\bm{P}}\left[\max_{\|\bm{\Delta}\|_{p}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})\right]=\mathbb{E}_{\bm{P}}[\bm{x}^{\top}C\bm{x}]+2\delta\mathbb{E}_{\bm{P}}\left[\|C\bm{x}\|_{q}\right]+\delta^{2}\mathbb{E}_{\bm{P}}\left[\bm{v}_{C\bm{x}}^{\top}C\bm{v}_{C\bm{x}}\right]+o(\delta^{2}),$
as $\delta\to 0$, where
$\bm{v}_{C\bm{x}}=\frac{\operatorname{sign}(C\bm{x})\odot|C\bm{x}|^{q-1}}{\|C\bm{x}\|_{q}^{q-1}}.$ (10)
Remark 4.2.
Since the term $\log\det C$ in the proposed framework (2) is deterministic, Proposition 4.1 focuses on the inner maximization problem in (2) to analyze the induced regularization under $\ell_{p}$-perturbation. Instead of only analyzing on the empirical distribution, we consider a more general distribution $\bm{P}$ satisfying certain regularity conditions. The stated regularity conditions in Proposition 4.1 ensure well-posedness and integrability.
Example 4.3.
We provide the examples for the cases when $p=2$ and $p=\infty$ in Proposition 4.1:
$\mathbb{E}_{\bm{P}}\left[\max_{\|\bm{\Delta}\|_{\infty}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})\right]$
$=$ $\mathbb{E}_{\bm{P}}[\bm{x}^{\top}C\bm{x}]+2\delta\mathbb{E}_{\bm{P}}\left[\|C\bm{x}\|_{1}\right]+\delta^{2}\mathbb{E}_{\bm{P}}\left[\operatorname{sign}(C\bm{x})^{\top}C\operatorname{sign}(C\bm{x})\right]+o(\delta^{2}),$
$\mathbb{E}_{\bm{P}}\left[\max_{\|\bm{\Delta}\|_{2}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})\right]$
$=$ $\mathbb{E}_{\bm{P}}[\bm{x}^{\top}C\bm{x}]+2\delta\mathbb{E}_{\bm{P}}\left[\|C\bm{x}\|_{2}\right]+\delta^{2}\mathbb{E}_{\bm{P}}\left[\frac{\bm{x}C^{3}\bm{x}}{\bm{x}C^{2}\bm{x}}\right]+o(\delta^{2}),$
as $\delta\to 0$.
Proposition 4.1 reveals that the proposed adversarially robust precision matrix framework is equivalent to the regularized precision matrix estimation in the asymptotic sense. For small perturbations, the worst-case loss inflates by a first-order term $2\delta\mathbb{E}_{\bm{P}}[\|C\bm{x}\|_{q}]$, a second-order term $\delta^{2}\mathbb{E}_{\bm{P}}[\bm{v}_{C\bm{x}}^{\top}C\bm{v}_{C\bm{x}}]$, and a higher-order residual.
The first-order regularization term arises from Hölder’s inequality and penalizes the $\ell_{q}$ norm of the gradient direction, where $1/p+1/q=1$. The vector $\bm{v}_{C\bm{x}}$ in the second-order term is the $\ell_{p}$-unit vector that achieves the associated Hölder dual pairing, as defined in (10). The second-order regularization term reflects the local curvature of the worst-case loss. Specifically, this term corresponds to the mean of the Rayleigh quotient of $C$ evaluated with respect to the vector $C\bm{x}$ (when $p=2$) or $\operatorname{sign}(C\bm{x})$ (when $p=\infty$). If $C$ has the eigen-decomposition $C=\sum_{j}\lambda_{j}\bm{u}_{j}\bm{u}_{j}^{\top}$, the associated Rayleigh quotient represents a weighted average of the eigenvalues. For $p=2$, it takes the form:
$\frac{\bm{x}^{\top}C^{3}\bm{x}}{\bm{x}^{\top}C^{2}\bm{x}}=\sum_{j}\lambda_{j}\left(\frac{\lambda_{j}^{2}(\bm{u}_{j}^{\top}\bm{x})^{2}}{\sum_{k}\lambda_{k}^{2}(\bm{u}_{k}^{\top}\bm{x})^{2}}\right).$ (11)
Similarly, for $p=\infty$:
$\operatorname{sign}(C\bm{x})^{\top}C\operatorname{sign}(C\bm{x})=\sum_{j}\lambda_{j}(\bm{u}_{j}^{\top}\operatorname{sign}(C\bm{x}))^{2}.$
Thus, the second-order regularization implies that adversarial perturbation induces a shrinkage effect on the eigenvalues of $C$. Furthermore, as seen in (11) for the $p=2$ case, this shrinkage effect is stronger on larger eigenvalues due to the $\lambda^{2}$ weighting.
We could establish the asymptotic distribution of the adversarially perturbed precision estimation utilizing the regularization effect in Proposition 4.1.
5 Asymptotic Distribution
This section will provide the asymptotic distribution of the adversarially perturbed estimator under different settings.
Recall that the proposed adversarially perturbed estimation is the solution to the following optimization problem:
$\widehat{C}\in\arg\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\max_{\|\bm{\Delta}\|_{p}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})\right]\right\}.$ (12)
Regarding the perturbation magnitude, we consider a regime where the perturbation level $\delta$ is small and decays as the sample size $n$ increases. Formally, we let $\delta=\delta_{n}=\eta n^{-\gamma}$ with constants $\eta,\gamma>0$.
The asymptotic distribution of the perturbed estimation $\widehat{C}$ admits an explicit expression as detailed below.
Theorem 5.1 (Asymptotic Distribution).
For problem (12), suppose $p\in(1,\infty]$, and $\delta_{n}=\eta n^{-\gamma},\gamma,\eta>0$, then we have that
  • If $0<\gamma<1/2$,
    $n^{\gamma}\left(\widehat{C}-\Sigma^{-1}\right)\Rightarrow-2\eta\Sigma^{-1}\mathbb{E}\left[\frac{\left(\mathrm{sign}(\Sigma^{-1}\bm{x})\odot|\Sigma^{-1}\bm{x}|^{q-1}\right)\bm{x}^{\top}}{\|\Sigma^{-1}\bm{x}\|_{q}^{q-1}}\right]\Sigma^{-1};$
  • If $\gamma=1/2$,
    $\sqrt{n}\left(\widehat{C}-\Sigma^{-1}\right)\Rightarrow\Sigma^{-1}\left(\mathbf{G}-2\eta\mathbb{E}\left[\frac{\left(\mathrm{sign}(\Sigma^{-1}\bm{x})\odot|\Sigma^{-1}\bm{x}|^{q-1}\right)\bm{x}^{\top}}{\|\Sigma^{-1}\bm{x}\|_{q}^{q-1}}\right]\right)\Sigma^{-1};$
  • If $\gamma>1/2$,
    $\sqrt{n}\left(\widehat{C}-\Sigma^{-1}\right)\Rightarrow\Sigma^{-1}\mathbf{G}\Sigma^{-1},$
where $\bm{x}\sim\mathcal{N}(\bm{0},\Sigma)$, $\mathrm{vec}(\mathbf{G})\sim\mathcal{N}(\bm{0},\Lambda)$, and $\Lambda$ is such that
$\mathrm{cov}([\mathbf{G}]_{ij},[\mathbf{G}]_{i^{\prime}j^{\prime}})=\mathrm{cov}([\bm{x}]_{i}[\bm{x}]_{j},[\bm{x}]_{i^{\prime}}[\bm{x}]_{j^{\prime}}).$
The results in Theorem 5.1 have some implications.
Firstly, the proposed perturbed estimation $\widehat{C}$ is a consistent estimator of the precision matrix $\Sigma^{-1}$, i.e., $\widehat{C}\to_{p}\Sigma^{-1}$, when the perturbation magnitude shrinks as the sample size increases. Then, the proposed perturbed estimation has a convergence order of $\sqrt{n}$ when $\gamma\geq 1/2$ and a slower convergence order of $n^{\gamma}$ when $\gamma\in(0,1/2)$.
Secondly, when we focus on cases where $\gamma\geq 1/2$, the perturbed estimation converges in distribution to a normal distribution.
Thirdly, the perturbation magnitude decays fast enough that the bias induced by the adversarial perturbation becomes asymptotically negligible when $\gamma>1/2$. In this regime, the estimation $\widehat{C}$ converges to a zero-mean normal distribution identical to that of the unperturbed inverse sample covariance, i.e., the maximum likelihood estimation from problem (1). This establishes a theoretical threshold: while adversarial training typically introduces a trade-off between robustness and accuracy (Tsipras et al., 2019), accuracy and perturbation are not in conflict when the perturbation order exceeds $1/2$ in our proposed adversarially perturbed precision matrix estimation framework.
Recall that we focus on the surrogate objective function instead when we discuss the $\ell_{\infty}$-perturbed estimation in Section 3, as shown in the following:
$\widetilde{C}\in\arg\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}C\bm{x}\right]+2\delta\sum_{j=1}^{d}\sum_{k=1}^{d}\omega_{k}|[C]_{kj}|+\delta^{2}\|C\|_{1,1}\right\}.$ (13)
In contrast to the asymptotic results in Theorem 5.2, the asymptotic distribution of the estimation $\widetilde{C}$ in problem (13) does not admit a simple closed form but can be characterized as the minimizer of a stochastic process, as shown in the following theorem.
Theorem 5.2 (Asymptotic Distribution when $p=\infty$ with Surrogate Objective).
If $\delta_{n}=\eta n^{-\gamma},\eta,\gamma>0$, for the estimation $\widetilde{C}$ obtained from problem (13), we have that
$\sqrt{n}\left(\widetilde{C}-\Sigma^{-1}\right)\Rightarrow\arg\min_{\bm{U}}W(\bm{U}).$
  • If $0<\gamma<1/2$,
    $W(\bm{U})=$ $\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma)$
    $+\sqrt{\frac{8}{\pi}}\eta\sum_{j=1}^{d}\sum_{k=1}^{d}[\Sigma]_{kk}\left([\bm{U}]_{jk}\operatorname{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)\right$.
    $\left.+|[\bm{U}]_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
  • If $\gamma=1/2$,
    $W(\bm{U})=$ $-\mathrm{tr}(\bm{U}\mathbf{G})+\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma)$
    $+\sqrt{\frac{8}{\pi}}\eta\sum_{j=1}^{d}\sum_{k=1}^{d}[\Sigma]_{kk}\left([\bm{U}]_{jk}\operatorname{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)\right$.
    $\left.+|[\bm{U}]_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
  • If $\gamma>1/2$,
    $W(\bm{U})=-\mathrm{tr}(\bm{U}\mathbf{G})+\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma),$
    i.e.,
    $\sqrt{n}\left(\widetilde{C}-\Sigma^{-1}\right)\Rightarrow\Sigma^{-1}\mathbf{G}\Sigma^{-1},$
where $\mathrm{vec}(\mathbf{G})\sim\mathcal{N}(\bm{0},\Lambda)$, and $\Lambda$ is such that
$\mathrm{cov}([\mathbf{G}]_{ij},[\mathbf{G}]_{i^{\prime}j^{\prime}})=\mathrm{cov}([\bm{x}]_{i}[\bm{x}]_{j},[\bm{x}]_{i^{\prime}}[\bm{x}]_{j^{\prime}}).$
Remark 5.3.
Theorem 5.2 provides the theoretical basis for establishing the sparsity recovery property of the estimation $\widetilde{C}$, as formalized in Corollary 3.2
Theorem 5.2 shows that the estimation $\widetilde{C}$ in problem (13) shares the same asymptotic normal distribution of the estimation $\widehat{C}$ in problem (12). However, for $\gamma=1/2$, the asymptotic distribution of $\widetilde{C}$ becomes non-normal. Similar to the standard $\ell_{1}$-regularized precision matrix estimation (Yuan and Lin, 2007), this non-normality helps encourage sparsity in some sense. Note that the asymptotic term $\Sigma]_{kk}$ presented in Theorem 5.2 represents the impact of adversarial perturbation in our surrogate framework (13).
To have a comprehensive understanding and comparison, we conclude some of the key characteristics of the asymptotic distribution of the perturbed estimation in Table 1.
$\in(1,\infty] p=p$$\infty$, surrogate Rate
$\gamma\in(0,1/2)$ constant constant $n^{\gamma}$
$\gamma=1/2$ normality with bias non-normality with bias, sparsity $\sqrt{n}$
$\gamma>1/2$ normality without bias $\sqrt{n}$
Table 1: Asymptotic Distribution of the Perturbed Estimation
6 Numerical Experiments
In this section, we will deliver numerical experiments for the proposed adversarially perturbed precision matrix estimation scheme and compare it with the $\ell_{1}$-regularized precision matrix estimation (Friedman et al., 2008). The experiments are implemented by R.
6.1 Computation
We will focus on the surrogate $\ell_{\infty}$-perturbed matrix estimation framework (9) discussed in Section 3.
Due to the symmetry of the matrix $C$, we can reformulate problem (9) equivalently as follows,
$\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathrm{tr}(\bar{A}C)+\|C\|_{1,\Lambda}\right\}$ (14)
where
$\bar{A}=\mathbb{E}_{\mathbb{P}_{n}}[\bm{x}\bm{x}^{\top}],\quad\|C\|_{1,\Lambda}=\sum_{k=1}^{d}\sum_{j=1}^{d}[\lambda]_{kj}\left|[C]_{kj}\right|$
$\lambda]_{kj}=\delta\mathbb{E}_{\mathbb{P}_{n}}\left[\left|[\bm{x}]_{k}\right|\right]+\delta\mathbb{E}_{\mathbb{P}_{n}}\left[\left|[\bm{x}]_{j}\right|\right]+\delta^{2}.$ (15)
The notation $\|C\|_{1,\Lambda}$ denotes the weighted element-wise $\ell_{1}$-norm of the matrix $C$ with the symmetric weight matrix $\Lambda$. The components of the weight matrix $\Lambda$ are dependent on the moment of the observations and the perturbation magnitude as shown in (15). In this way, problem (14) can be considered as the generalized sparse inverse covariance estimation problem discussed in Hsieh et al. (2014). Thus, problem (14) can be solved by the QUIC algorithm proposed by Hsieh et al. (2014). The QUIC algorithm is a second-order algorithm and has been proven to be superlinearly convergent (Hsieh et al., 2014).
In addition, immediately from Theorem 9 in Hsieh et al. (2014), we can conclude that the precision matrix estimation obtained from the $\ell_{\infty}$-perturbed precision matrix estimation proposed in Section 3 is unique.
Corollary 6.1.
There exists a unique minimizer $\widetilde{C}$ for the optimization problem (14) when $\delta>0$.
As the $\ell_{\infty}$-perturbed precision matrix estimation has the weight matrix dependent on the observed samples, the $\ell_{1}$-regularized precision matrix estimation (Friedman et al., 2008) has a constant weight, as shown in the following:
$\inf_{C\succ 0}\left\{-\log\mathrm{det}C+\mathrm{tr}(\bar{A}C)+\lambda\|C\|_{1}\right\}.$ (16)
We also apply the QUIC algorithm (Hsieh et al., 2014) to solve the optimization problem above.
In practice, we do not penalize the diagonals for either method to promote more desirable numerical performance for all experiments in this section.
6.2 Experiments with Synthetic Data
In this subsection, we will carry out the numerical experiments on the following synthetic models, as suggested in Yuan and Lin (2007).
Model 1. An $AR(1)$ model with $c_{i,i}=1$ and $c_{i,i-1}=c_{i-1,i}=0.5$, for i=1,,di=1,...,d.
Model 2. An $AR(2)$ model with $c_{i,i}=1$, $c_{i,i-1}=c_{i-1,i}=0.5$ and $c_{i,i-2}=c_{i-2,i}=0.25$, for i=1,,di=1,...,d.
Model 3. An $AR(3)$ model with $c_{i,i}=1$, $c_{i,i-1}=c_{i-1,i}=0.4$ and $c_{i,i-2}=c_{i-2,i}=c_{i,i-3}=c_{i-3,i}=0.2,$ for i=1,,di=1,...,d.
Model 4. An $AR(4)$ model with $c_{i,i}=1$, $c_{i,i-1}=c_{i-1,i}=0.4$, $c_{i,i-2}=c_{i-2,i}=c_{i,i-3}=c_{i-3,i}=0.2$ and $c_{i,i-4}=c_{i-4,i}=0.1,$ for i=1,,di=1,...,d.
Model 5. Star model with every node connected to the first node, with $c_{i,i}=1$, $c_{1,i}=c_{i,1}=0.2$ and $c_{i,j}=0$ otherwise, for i,j=1,,di,j=1,...,d.
Model 6. Circle model with $c_{ii}=1$, $c_{i,i-1}=c_{i-1,i}=0.5$ and $c_{1,d}=c_{d,1}=0.4$, for i=1,,di=1,...,d.
We may use the following metrics to evaluate the numerical performance in the rest of this section: true negative rate (TNR), true positive rate (TPR), accuracy (ACC), and Matthews Correlation Coefficient (MCC) (Guilford, 1954). The metrics are defined as follows:
$\mathrm{TNR}=\frac{\mathrm{TN}}{\mathrm{TN}+\mathrm{FP}},\quad\mathrm{TPR}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}},$
$\mathrm{ACC}=\frac{\mathrm{TN}+\mathrm{TP}}{\mathrm{TN}+\mathrm{FP}+\mathrm{TP}+\mathrm{FN}}$
$\mathrm{MCC}=\frac{\mathrm{TP}\times\mathrm{TN}-\mathrm{FP}\times\mathrm{FN}}{\sqrt{(\mathrm{TP}+\mathrm{FP})(\mathrm{TP}+\mathrm{FN})(\mathrm{TN}+\text{FP})(\text{TN}+\text{FN})}},$
where $\mathrm{TP}$, $\mathrm{TN}$, $\mathrm{FP}$, $\mathrm{FN}$ denote the number of true positives, true negatives, false positives, and false negatives, respectively.
For each model, we run simulations by the QUIC algorithm (Hsieh et al., 2014) with dimension $d=20$ and sample sizes $n=\{30,50,70\}$. We consider an equally spaced grid for the perturbation magnitude $\delta$ on $[0.01,0.3]$ in problem (14) and select the regularization parameter $\lambda$ in problem (16) over an equally spaced grid on $[0.01,0.8]$. For each setting, we run two methods 50 times and record the average and standard error.
6.2.1 Sparsity Recovery and Accuracy
We first compare the performance, i.e., sparsity recovery ability and accuracy, based on Kullback–Leibler (KL) loss, TNR, and MCC. The KL loss is defined as follows:
$\mathrm{KL}=-\log|\bar{C}|+\mathrm{tr}(\bar{C}\Sigma)-\left(-\log|\Sigma^{-1}|+d\right),$
where $\Sigma$ is the ground-truth covariance matrix, and $\bar{C}$ is the estimation of the precision matrix.
We visualize the experimental results for the 6 models in Figure 1, Figure 2, and Figure 3. In the figures, “Perturbed” refers to the performance of our proposed method, whereas “Standard” indicates the performance of the standard $\ell_{1}$-regularized method. The x-axis shows the values of the $\delta$ and, $\lambda$ while the y-axis shows the different metrics.
According to the figures, we could conclude that the proposed method (14) has the ability of sparsity recovery since the TNR of the perturbed method can reach or be close to $1$. In addition, the performances of two methods are comparable in terms of accuracy and sparsity recovery ability.
Figure 1: Accuracy and Sparsity Recovery on $AR(1)$ and $AR(2)$
Figure 2: Accuracy and Sparsity Recovery on $AR(3)$ and $AR(4)$
Figure 3: Accuracy and Sparsity Recovery on the Circle model and the Star model
6.2.2 Adversarial Robustness
In addition to the sparsity recovery ability and the accuracy discussed previously, we also evaluate the adversarial robustness of the methods.
In each simulation trial, three independent datasets are generated from the true distribution: a training set $X_{\text{train}}$ of size $n$, a validation set of size $n$, and a test set $X_{\text{test}}$ of size $5n$. For each sample size $n$, we first identify the optimal choice of $\delta$ for the proposed method and $\lambda$ for the standard $\ell_{1}$-regularized method by minimizing the value of $\mathrm{KL}$ on the validation set. We then derive the estimation $\bar{C}_{1}$ from the training data. Next, we generate adversarial examples from the test data based on $\bar{C}_{1}$. The adversarial perturbation is constructed using a gradient-sign method defined as:
$X_{\text{adv}}=X_{\text{test}}+\varepsilon\text{sign}(X_{\text{test}}\bar{C}_{1}),$
where $\varepsilon$ controls the magnitude of the adversarial perturbation. We vary $\varepsilon$ across a grid from 0 to $1$.
Then, we run the two estimation methods based on the perturbed data $X_{\text{adv}}$ and obtain the estimation $\bar{C}_{2}$. The robustness of each method is evaluated by the $\mathrm{KL}$ loss of $\bar{C}_{2}$.
$\text{adversarial KL}=-\log|\bar{C}_{2}|+\mathrm{tr}(\bar{C}_{2}\Sigma)-\left(-\log|\Sigma^{-1}|+d\right)$
A smaller adversarial KL loss indicates that the estimator is more robust to adversarial data perturbation.
The results are visualized in Figure 4, Figure 5, and Figure 6. The x-axis shows the values of $\varepsilon$ while the value of the adversarial KL. The figures demonstrate that the proposed perturbed method is more adversarially robust than the standard $\ell_{1}$-regularized model for most of the scenarios.
Figure 4: Adversarial Robustness on $AR(1)$ and $AR(2)$
Figure 5: Adversarial Robustness on $AR(3)$ and $AR(4)$
Figure 6: Adversarial Robustness on the Circle model and the Star model
6.3 Experiments with Real Data
In this subsection, we run the two estimation methods on the leukemia dataset (Golub et al., 1999) to compare the associated real-world practical performance.
The dataset comprises 7129 gene expression profiles collected from 72 patients, of whom 47 were diagnosed with Acute Lymphoblastic Leukemia (ALL) and 25 with Acute Myeloid Leukemia (AML). We first select the top $d$ genes with the highest marginal variance. The dimension $d$ is varied across the set $\{20,30,40,50,60\}$. The data is centered and scaled to unit variance within the training set and test set, respectively. Following the Linear Discriminant Analysis (LDA) framework, the classification rule relies on the LDA score. For the observation $\bm{x}$, the LDA score is defined by
$\delta_{k}(\bm{x})=\bm{x}^{T}\hat{C}\hat{\bm{\mu}}_{k}-\frac{1}{2}\hat{\bm{\mu}}_{k}^{T}\hat{C}\hat{\bm{\mu}}_{k}+\log\hat{\pi}_{k},\quad k=1,2,$
where $\hat{\pi}_{k}$ represents the proportion of group $k$ in the training set and $\hat{\bm{\mu}}_{k}$ is the empirical mean of the observations in the training set. The classification rule is $\hat{k}(\bm{x})=\arg\max_{k\in\{1,2\}}\delta_{k}(\bm{x})$. The classification performance is dependent on the estimation accuracy of the precision matrix.
The experiment employs a nested cross-validation scheme. In the outer loop, the data is split into 10 stratified folds. For each training run, we additionally carry out an inner stratified 5‑fold cross‑validation to tune the parameters ($\delta$ in the perturbed method (14) and $\lambda$ in the standard $\ell_{1}$-regularized method (16)) through a grid search within $[0.01,1]$, choosing the value that has the best classification accuracy. With the optimal parameter, we train the model on the full training set and then evaluate the model performance on the held‑out test fold. We replicate this whole scheme $50$ times.
To evaluate the performance, we use ACC, MCC, TNR, and TPR. The average value and the standard error of the metrics for 50 replications are recorded in Table 2. Notably, our perturbed method yields consistently higher ACC and MCC, at the cost of a slight decrease in TPR. In conclusion, the proposed method demonstrates superior efficacy in the precision matrix estimation for the leukemia dataset, outperforming the standard $\ell_{1}$-regularized method.
Method ACC MCC TNR TPR
$d=20$ Perturbed 0.868(0.00360) 0.691 (0.01009) 0.627 (0.01014) 0.995 (0.00137)
Standard 0.864 (0.00313) 0.679 (0.00882) 0.611 (0.00951) 0.997 (0.00101)
$d=30$ Perturbed 0.910 (0.00290) 0.792 (0.00806) 0.741(0.00886) 1.000 (0.000500)
Standard 0.902 (0.00301) 0.775 (0.00760) 0.716 (0.00871) 1.000 (0.000000)
$d=40$ Perturbed 0.933 (0.00285) 0.850 (0.00732) 0.809 (0.00840) 1.000 (0.00040)
Standard 0.921 (0.00305) 0.819 (0.00807) 0.771 (0.00928) 1.000 (0.00000)
$d=50$ Perturbed 0.927 (0.00312) 0.833 (0.00841) 0.798 (0.00926) 0.994 (0.00153)
Standard 0.916 (0.00268) 0.808 (0.00722) 0.760 (0.00778) 0.998 (0.000928)
$d=60$ Perturbed 0.929(0.00247) 0.839 (0.00712) 0.806 (0.00727) 0.993 (0.00143)
Standard 0.912 (0.00305) 0.802 (0.00779) 0.757 (0.00800) 0.994 (0.00148)
Table 2: Average Value and Standard Error of the Metrics
7 Discussion
This paper introduces a versatile perturbed estimation framework for the precision matrix and derives the associated regularization effects and asymptotic distributions. Among different perturbation geometries, the $\ell_{2}$- and $\ell_{\infty}$-perturbations are specifically analyzed. Notably, we demonstrate that the proposed framework recovers Wasserstein shrinkage estimation, achieves sparsity recovery, and improves adversarial robustness. In addition to these properties, future directions may include extending our framework to handle complex data structures, including the unknown group structures studied in (Cheng et al., 2025) and the compositional data investigated in (Zhang et al., 2025).
Appendix A Proofs
A.1 Proof of Theorem 2.1
The proof for Theorem 2.1 is based on the following lemma.
Lemma A.1.
The dual of the optimization problem
$\max_{\|\bm{\Delta}\|_{2}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})$
is the following optimization problem
$\bm{x}^{\top}C\bm{x}+\inf_{\lambda I-C\succ 0}\left\{\lambda\delta^{2}+\bm{x}^{\top}C(\lambda I-C)^{-1}C\bm{x}\right\}.$
Proof.
The Lagrangian dual of $\max_{\|\bm{\Delta}\|_{2}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})$ is as follows:
$\inf_{\lambda\geq 0}\left\{\lambda\delta^{2}+\max_{\bm{\Delta}}\left\{(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})-\lambda\|\bm{\Delta}\|_{2}^{2}\right\}\right\},$
which is equivalent to
$\bm{x}^{\top}C\bm{x}+\inf_{\lambda\geq 0}\left\{\lambda\delta^{2}+\max_{\bm{\Delta}}\left\{\bm{\Delta}^{\top}(C-\lambda I)\bm{\Delta}+2\bm{x}^{\top}C\bm{\Delta}\right\}\right\}.$
To make the inner optimization problem finite, we should require $\lambda I-C\succ 0$. In this way, the optimal value will be obtained when
$\bm{\Delta}=(\lambda I-C)^{-1}C\bm{x},$
the resulting optimal value is
$\max_{\bm{\Delta}\in\mathbb{R}^{p}}\left\{\bm{\Delta}^{\top}(C-\lambda I)\bm{\Delta}+2\bm{x}^{\top}C\bm{\Delta}\right\}=\bm{x}^{\top}C(\lambda I-C)^{-1}C\bm{x}$
Now we begin to prove Theorem 2.1.
Proof.
It follows from Lemma A.1 that problem (3) is equivalent to the following:
$\inf_{C\succ 0,\lambda I-C\succ 0}\left\{-\log\mathrm{det}C+\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}^{\top}C\bm{x}_{i}+\lambda\delta^{2}+\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}^{\top}C(\lambda I-C)^{-1}C\bm{x}_{i}\right\}.$
By introducing new variables $t_{i}$, we have that
$\inf_{C\succ 0,\lambda I-C\succ 0}\left\{-\log\mathrm{det}C+\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}^{\top}C\bm{x}_{i}+\lambda\delta^{2}+\frac{1}{n}\sum_{i=1}^{n}t_{i}\right\}.$
$\mathrm{s.t.}\quad\bm{x}_{i}^{\top}C(\lambda I-C)^{-1}C\bm{x}_{i}\leq t_{i}\quad\text{for }1\leq i\leq n.$
Due to the Schur complement, we have the following reformulation:
$\inf_{\lambda,\{t_{i}\},C}$ $-\log\mathrm{det}C+\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}^{\top}C\bm{x}_{i}+\lambda\delta^{2}+\frac{1}{n}\sum_{i=1}^{n}t_{i}$
subject to $C\succ 0,\lambda I-C\succ 0$
$\begin{pmatrix}t_{i}&\bm{x}_{i}^{\top}C\\ C\bm{x}_{i}&\lambda I-C\end{pmatrix}\succeq 0,\quad\forall i=1,\dots,n.$
A.2 Proof of Theorem 2.3
Proof.
It remains to show the following identity:
$\mathrm{tr}(C(\lambda I-C)^{-1}C\bar{A})=\lambda^{2}\mathrm{tr}((\lambda I-C)^{-1}\bar{A})-\mathrm{tr}((\lambda I+C)\bar{A}).$
Since a matrix can commute with its own resolvent matrix, we have that
$\mathrm{tr}(C(\lambda I-C)^{-1}C\bar{A})=\mathrm{tr}(C^{2}(\lambda I-C)^{-1}\bar{A}).$
Then, we can conclude that
$\mathrm{tr}(C(\lambda I-C)^{-1}C\bar{A})-\lambda^{2}\mathrm{tr}((\lambda I-C)^{-1}\bar{A})=\mathrm{tr}((C^{2}-\lambda^{2}I)(\lambda I-C)^{-1}\bar{A}).$
Since
$(C^{2}-\lambda^{2}I)=(C-\lambda I)(C+\lambda I),$
the identity has been proven. ∎
A.3 Proof of Proposition 3.1
Proof.
When $p\geq 1$, we have that
$\max_{\|\bm{\Delta}\|_{p}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})$ $\leq\max_{\|\bm{\Delta}\|_{\infty}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})$
$=\bm{x}^{\top}C\bm{x}+\max_{\|\bm{\Delta}\|_{\infty}\leq\delta}\left\{2\bm{x}^{\top}C\bm{\Delta}+\bm{\Delta}^{\top}C\bm{\Delta}\right\}$
$\leq\bm{x}^{\top}C\bm{x}+\max_{\|\bm{\Delta}\|_{\infty}\leq\delta}2\bm{x}^{\top}C\bm{\Delta}+\max_{\|\bm{\Delta}\|_{\infty}\leq\delta}\bm{\Delta}^{\top}C\bm{\Delta}$
$=\bm{x}^{\top}C\bm{x}+2\delta\|C\bm{x}\|_{1}+\max_{\|\bm{\Delta}\|_{\infty}\leq\delta}\bm{\Delta}^{\top}C\bm{\Delta}$
$\leq\bm{x}^{\top}C\bm{x}+2\delta\|C\bm{x}\|_{1}+\delta^{2}\|C\|_{1,1}$
Notice that
$\mathbb{E}_{\mathbb{P}_{n}}\left[\|C\bm{x}\|_{1}\right]=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{d}|[C]_{j}\bm{x}_{i}|$ $\leq\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{d}\sum_{k=1}^{d}|[C]_{kj}||[\bm{x}_{i}]_{k}|$
$=\sum_{j=1}^{d}\sum_{k=1}^{d}|[C]_{kj}|\left(\frac{1}{n}\sum_{i=1}^{n}|[\bm{x}_{i}]_{k}|\right).$
We can conclude that the inequality holds. ∎
A.4 Proof of Proposition 4.1
We first state two propositions before we prove Proposition 4.1.
Proposition A.2.
For vector $b\in\mathbb{R}^{d}$, positive definite matrix $Q\in\mathbb{R}^{d\times d}$ and $p\in(1,\infty)$, we have that
$\max_{\|\Delta\|_{p}\leq 1}b^{\top}\Delta+\varepsilon\Delta^{\top}Q\Delta=\|b\|_{q}+\varepsilon\widetilde{\Delta}^{\top}Q\widetilde{\Delta}+o(\varepsilon),$ (17)
as $\varepsilon\to 0$, where $1/p+1/q=1$, and $\widetilde{\Delta}$ is the unique solution to the problem
$\max_{\|\Delta\|_{p}\leq 1}b^{\top}\Delta.$
Proof.
We first prove that the optimization problem (17) has a unique solution when $\varepsilon\to 0$.
The Lagrangian function of problem (17) is
$L_{\varepsilon}(\Delta,\lambda)=b^{\top}\Delta+\varepsilon\Delta^{\top}Q\Delta-\lambda(\|\Delta\|_{p}^{p}-1).$
Since the constraint set $\|\Delta\|_{p}\leq 1$ is convex, it follows from the KKT condition that the solution to the problem, denoted by $\Delta_{\ast}$ and $\lambda_{\ast}$, follows the equations:
$\|\Delta_{\ast}\|_{p}^{p}-1=0,\quad b+\varepsilon 2Q\Delta_{\ast}-\lambda_{\ast}p\|\Delta_{\ast}\|_{p}^{p-1}|\Delta_{\ast}|^{p-1}\odot\operatorname{sign}(\Delta_{\ast})=0.$ (18)
Assume there are two distinct solutions to problem (17) when $\varepsilon\to 0$, denoted by $\Delta_{\ast}^{1}$ and $\Delta_{\ast}^{2}$. Equivalently, both $\Delta_{\ast}^{1}$ and $\Delta_{\ast}^{2}$ satisfy the equations in (18), indicating
$2\varepsilon Q(\Delta_{\ast}^{1}-\Delta_{\ast}^{2})=\lambda_{\ast}p(|\Delta_{\ast}^{2}|^{p-1}\odot\operatorname{sign}(\Delta_{\ast}^{2})-|\Delta_{\ast}^{1}|^{p-1}\odot\operatorname{sign}(\Delta_{\ast}^{1})),$ (19)
$\|\Delta_{\ast}^{1}\|_{p}=\|\Delta_{\ast}^{2}\|_{p}=1,\quad\Delta_{\ast}^{1}\not=\Delta_{\ast}^{2}$
It is easy to see that the left-hand side of (19) goes to $\bm{0}$ when $\varepsilon\to 0$ while the right-hand side of (19) strictly not equal to $\bm{0}$ when $p\in(1,\infty)$ and $\Delta_{\ast}^{1}\not=\Delta_{\ast}^{2}$. This produces a contradiction. Thus, the optimization problem (17) has a unique solution when $\varepsilon\to 0$ and $p\in(1,\infty)$.
Then, since the objective function is convex, continuous, and differentiable w.r.t $\varepsilon$ for fixed $\Delta$, by Danskin’s theorem (Bonnans and Shapiro, 2013), if we denote
$\phi_{b,Q}(\varepsilon)=\max_{\|\Delta\|_{p}\leq 1}b^{\top}\Delta+\varepsilon\Delta^{\top}Q\Delta,\quad\Delta_{b,Q}(\varepsilon)=\arg\max_{\|\Delta\|_{p}\leq 1}b^{\top}\Delta+\varepsilon\Delta^{\top}Q\Delta,$
then we have that
$\frac{d}{d\varepsilon}\phi_{b,Q}(\varepsilon)=\Delta_{b,Q}(\varepsilon)^{\top}Q\Delta_{{b,Q}}(\varepsilon),$
and
$\frac{d}{d\varepsilon}\phi_{b,Q}(\varepsilon)\Big|_{\varepsilon=0}=\widetilde{\Delta}^{\top}Q\widetilde{\Delta}$,
It follows from Taylor’s expansion that
$\phi_{b,Q}(\varepsilon)=\phi_{b,Q}(0)+\varepsilon\widetilde{\Delta}^{\top}Q\widetilde{\Delta}+o(\varepsilon)=\|b\|_{q}+\varepsilon\widetilde{\Delta}^{\top}Q\widetilde{\Delta}+o(\varepsilon),$
as $\varepsilon\to 0$. ∎
Proposition A.3.
For vector $b\in\mathbb{R}^{d}$ with $b]_{i}\not=0$ for i=1,,di=1,...,d, positive definite matrix $Q\in\mathbb{R}^{d\times d}$, we have that
$\max_{\|\Delta\|_{\infty}\leq 1}b^{\top}\Delta+\varepsilon\Delta^{\top}Q\Delta=\|b\|_{1}+\varepsilon\operatorname{sign}(b)^{\top}Q\operatorname{sign}(b),$ (20)
as $\varepsilon\to 0$.
Proof.
When $\varepsilon=0$, the optimization problem $\max_{\|\Delta\|_{\infty}\leq 1}b^{\top}\Delta$ has the unique solution at $\operatorname{sign}(b)$.
When $\varepsilon\not=0$, the objective function is convex and the constraint set is a convex compact set. In this way, the solution to the optimization problem is taken at the boundary of the constraint set, i.e., the solution $\Delta^{\ast}$ satisfies that $\Delta^{\ast}]_{i}=\{-1,1\}$. Since $b]_{i}\not=0$, when $\varepsilon$ is very small, flipping the sign of $\Delta^{\ast}$ will decrease the value of the objective function. Thus, the optimal solution is $\operatorname{sign}(b)$ when $\varepsilon\to 0$. ∎
Now we begin to prove Proposition 4.1.
Proof.
Notice we have that
$\max_{\|\bm{\Delta}\|_{p}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})=\bm{x}^{\top}C\bm{x}+\delta\max_{\|\bm{\Delta}\|_{p}\leq 1}\left\{2(C\bm{x})^{\top}\bm{\Delta}+\delta\bm{\Delta}^{\top}C\bm{\Delta}\right\}.$
In the probability space $(\bm{P},C,\mathcal{F})$ for variable $\bm{x}$, for each $\bm{\omega}\in C$, it follows from Proposition A.2 and Proposition A.3 that
$\max_{\|\bm{\Delta}\|_{p}\leq 1}\left\{2(C\bm{x}(\bm{\omega}))^{\top}\bm{\Delta}+\delta\bm{\Delta}^{\top}C\bm{\Delta}\right\}=2\|C\bm{x}(\bm{\omega})\|_{q}+\delta\bm{v}_{\bm{x}(\bm{\omega})}^{\top}C\bm{v}_{\bm{x}(\bm{\omega})}+\mathcal{R}(\bm{\omega}),$
where
$\lim_{\delta\to 0}\frac{\mathcal{R}(\bm{\omega})}{\delta}=0.$
Since we have that
$|\mathcal{R}(\bm{\omega})|$ $=\left|\max_{\|\bm{\Delta}\|_{p}\leq 1}\left\{2(C\bm{x}(\bm{\omega}))^{\top}\bm{\Delta}+\delta\bm{\Delta}^{\top}C\bm{\Delta}\right\}-2\|C\bm{x}(\bm{\omega})\|_{q}+\delta\bm{v}_{\bm{x}(\bm{\omega})}^{\top}C\bm{v}_{\bm{x}(\bm{\omega})}\right|$ (21)
$\leq 4\|C\bm{x}(\bm{\omega})\|_{q}+2\delta\max_{\|y\|_{p}\leq 1,y\in\mathbb{R}^{d}}y^{\top}Cy$
$\leq 4\|C\|_{p\to q}\|\bm{x}(\bm{\omega})\|_{p}+2\delta\|C\|_{p\to q},$
implying that $\mathcal{R}(\bm{\omega})$ is upper bounded by an integral function. Due to the dominated convergence theorem, we have that
$\lim_{\delta\to 0}\frac{\mathbb{E}_{\bm{P}}[\mathcal{R}]}{\delta}=\int_{\bm{\omega}\in C}\lim_{\delta\to 0}\frac{\mathcal{R}(\bm{\omega})}{\delta}d\bm{P}(\bm{\omega})=0.$
A.5 Proof of Theorem 5.1
Proof.
We define
$\Psi_{n}(C)=-\log\det(C)+\mathbb{E}_{\mathbb{P}_{n}}\left[\max_{\|\bm{\Delta}\|_{p}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C(\bm{x}+\bm{\Delta})\right].$
For $C\succ 0$, the function $\Psi_{n}(C)$ is convex since we have the following:
$\Psi_{n}(\alpha C_{1}+(1-\alpha)C_{2})$ (22)
$=$ $-\log\det(\alpha C_{1}+(1-\alpha)C_{2})+\mathbb{E}_{\mathbb{P}_{n}}\left[\max_{\|\bm{\Delta}\|_{p}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}(\alpha C_{1}+(1-\alpha)C_{2})(\bm{x}+\bm{\Delta})\right]$
$\leq$ $-\alpha\log\det(C_{1})-(1-\alpha)\log\det(C_{2})$
$+$ $\alpha\mathbb{E}_{\mathbb{P}_{n}}\left[\max_{\|\bm{\Delta}\|_{p}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C_{1}(\bm{x}+\bm{\Delta})\right]+(1-\alpha)\mathbb{E}_{\mathbb{P}_{n}}\left[\max_{\|\bm{\Delta}\|_{p}\leq\delta}(\bm{x}+\bm{\Delta})^{\top}C_{2}(\bm{x}+\bm{\Delta})\right],$
where the second inequality comes from the convexity of the function $\log\det(C)$.
Case 1: $\gamma=1/2$.
It follows from Proposition 4.1 that
$V_{n}(\bm{U})$
$\overset{\vartriangle}{=}$ $n\left(\Psi_{n}(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})-\Psi(\Sigma^{-1})\right)$
$=$ $-n\log\det(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})+n\log\det(\Sigma^{-1})+\sqrt{n}\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}\bm{U}\bm{x}\right]$
$+2\eta\sqrt{n}\left(\mathbb{E}_{\mathbb{P}_{n}}\left[\left\|(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})\bm{x}\right\|_{q}\right]-\mathbb{E}_{\mathbb{P}_{n}}\left[\left\|\Sigma^{-1}\bm{x}\right\|_{q}\right]\right)$
$+\eta^{2}\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{v}_{\Sigma^{-1}\bm{x}}^{\top}\Sigma^{-1}\bm{v}_{\Sigma^{-1}\bm{x}}\right]-\eta^{2}\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{v}_{(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})\bm{x}}^{\top}\left(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U}\right)\bm{v}_{(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})\bm{x}}\right]$
$+o(1).$
It follows from the proof of Theorem 1 in Yuan and Lin (2007) that
$\log\det(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})+\log\det(\Sigma^{-1})=-\frac{\mathrm{tr}(\bm{U}\Sigma)}{\sqrt{n}}+\frac{1}{2}\frac{\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma)}{n}+o\left(\frac{1}{n}\right).$
We can rewrite the term $\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}\bm{U}\bm{x}\right]$ as $\mathrm{tr}(\bm{U}\bar{A})$, where
$\bar{A}=\frac{1}{n}\sum_{i=1}^{n}\bm{x}_{i}\bm{x}_{i}^{\top}.$
In this way, we have that
$-n\log\det(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})+n\log\det(\Sigma^{-1})+\sqrt{n}\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}\bm{U}\bm{x}\right]$ (23)
$=$ $\mathrm{tr}(\bm{U}\sqrt{n}(\bar{A}-\Sigma))+\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma)+o(1)$
$\Rightarrow$ $-\mathrm{tr}(\bm{U}\mathbf{G})+\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma),$
where $\mathrm{vec}(\mathbf{G})\sim\mathcal{N}(\bm{0},\Lambda)$, and $\Lambda$ is such that
$\mathrm{cov}([\mathbf{G}]_{ij},[\mathbf{G}]_{i^{\prime}j^{\prime}})=\mathrm{cov}([\bm{x}]_{i}[\bm{x}]_{j},[\bm{x}]_{i^{\prime}}[\bm{x}]_{j^{\prime}}).$
Notice that we also have
$\sqrt{n}\left(\mathbb{E}_{\mathbb{P}_{n}}\left[\left\|(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})\bm{x}\right\|_{q}\right]-\mathbb{E}_{\mathbb{P}_{n}}\left[\left\|\Sigma^{-1}\bm{x}\right\|_{q}\right]\right)$ (24)
$\to_{p}\mathbb{E}\left[\frac{\left(\mathrm{sign}(\Sigma^{-1}\bm{x})\odot|\Sigma^{-1}\bm{x}|^{q-1}\right)^{\top}\bm{U}\bm{x}}{\|\Sigma^{-1}\bm{x}\|_{q}^{q-1}}\right].$
Thus,
$V_{n}(\bm{U})\Rightarrow V(\bm{U}),$
where
$V(\bm{U})=-\mathrm{tr}(\bm{U}\mathbf{G})+\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma)+2\eta\mathbb{E}\left[\frac{\left(\mathrm{sign}(\Sigma^{-1}\bm{x})\odot|\Sigma^{-1}\bm{x}|^{q-1}\right)^{\top}\bm{U}\bm{x}}{\|\Sigma^{-1}\bm{x}\|_{q}^{q-1}}\right].$
We take the gradient of $V(\bm{U})$ w.r.t. $\bm{U}$ and obtain that
$\mathbf{G}+\Sigma\bm{U}^{\ast}\Sigma+2\eta\mathbb{E}\left[\frac{\left(\mathrm{sign}(\Sigma^{-1}\bm{x})\odot|\Sigma^{-1}\bm{x}|^{q-1}\right)\bm{x}^{\top}}{\|\Sigma^{-1}\bm{x}\|_{q}^{q-1}}\right]=0,$
$\bm{U}^{\ast}=\Sigma^{-1}\left(\mathbf{G}-2\eta\mathbb{E}\left[\frac{\left(\mathrm{sign}(\Sigma^{-1}\bm{x})\odot|\Sigma^{-1}\bm{x}|^{q-1}\right)\bm{x}^{\top}}{\|\Sigma^{-1}\bm{x}\|_{q}^{q-1}}\right]\right)\Sigma^{-1}.$
Since $V(\bm{U})$ is strongly convex due to the positive definiteness of $\Sigma$, we have that
$\sqrt{n}\left(\widehat{C}-\Sigma^{-1}\right)\Rightarrow\Sigma^{-1}\left(\mathbf{G}-2\eta\mathbb{E}\left[\frac{\left(\mathrm{sign}(\Sigma^{-1}\bm{x})\odot|\Sigma^{-1}\bm{x}|^{q-1}\right)\bm{x}^{\top}}{\|\Sigma^{-1}\bm{x}\|_{q}^{q-1}}\right]\right)\Sigma^{-1}.$
Case 2: $\gamma>1/2$.
$V_{n}(\bm{U})$
$\overset{\vartriangle}{=}$ $n\left(\Psi_{n}(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})-\Psi(\Sigma^{-1})\right)$
$=$ $-n\log\det(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})+n\log\det(\Sigma^{-1})+\sqrt{n}\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}\bm{U}\bm{x}\right]$
$+2\eta n^{1/2-\gamma}\sqrt{n}\left(\mathbb{E}_{\mathbb{P}_{n}}\left[\left\|(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})\bm{x}\right\|_{q}\right]-\mathbb{E}_{\mathbb{P}_{n}}[\left\|\Sigma^{-1}\bm{x}\right\|_{q}]\right)$
$+\eta^{2}n^{1-2\gamma}\left(\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{v}_{\Sigma^{-1}\bm{x}}^{\top}\Sigma^{-1}\bm{v}_{\Sigma^{-1}\bm{x}}\right]-\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{v}_{(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})\bm{x}}^{\top}\left(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U}\right)\bm{v}_{(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})\bm{x}}\right]\right)$
$+o(1).$
Since $\gamma>1/2$, we have that
$V_{n}(\bm{U})\Rightarrow V(\bm{U}),$
where
$V(\bm{U})=-\mathrm{tr}(\bm{U}\mathbf{G})+\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma).$
Then, we have that
$\sqrt{n}\left(\widehat{C}-\Sigma^{-1}\right)\Rightarrow\Sigma^{-1}\mathbf{G}\Sigma^{-1}.$
Case 3: $0<\gamma<1/2$.
$V_{n}(\bm{U})\overset{\vartriangle}{=}$
$n^{2\gamma}\left(\Psi_{n}(\Sigma^{-1}+\frac{1}{n^{\gamma}}\bm{U})-\Psi(\Sigma^{-1})\right)$
$=$ $-n^{2\gamma}\log\det(\Sigma^{-1}+\frac{1}{n^{\gamma}}\bm{U})+n^{2\gamma}\log\det(\Sigma^{-1})+n^{\gamma}\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}\bm{U}\bm{x}\right]$
$+2\eta n^{\gamma}\left(\mathbb{E}_{\mathbb{P}_{n}}\left[\left\|(\Sigma^{-1}+\frac{1}{n^{\gamma}}\bm{U})\bm{x}\right\|_{q}\right]-\mathbb{E}_{\mathbb{P}_{n}}[\left\|\Sigma^{-1}\bm{x}\right\|_{q}]\right)$
$+\eta^{2}\left(\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{v}_{\Sigma^{-1}\bm{x}}^{\top}\Sigma^{-1}\bm{v}_{\Sigma^{-1}\bm{x}}\right]-\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{v}_{(\Sigma^{-1}+\frac{1}{n^{\gamma}}\bm{U})\bm{x}}^{\top}\left(\Sigma^{-1}+\frac{1}{n^{\gamma}}\bm{U}\right)\bm{v}_{(\Sigma^{-1}+\frac{1}{n^{\gamma}}\bm{U})\bm{x}}\right]\right)$
$+o(1).$
We have that
$n^{2\gamma}\log\det(\Sigma^{-1}+\frac{1}{n^{\gamma}}\bm{U})+n^{2\gamma}\log\det(\Sigma^{-1})+n^{\gamma}\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}\bm{U}\bm{x}\right]\Rightarrow\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma).$
We conclude that
$V_{n}(\bm{U})\Rightarrow V(\bm{U}),$
where
$V(\bm{U})=\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma)+2\eta\mathbb{E}\left[\frac{\left(\mathrm{sign}(\Sigma^{-1}\bm{x})\odot|\Sigma^{-1}\bm{x}|^{q-1}\right)^{\top}\bm{U}\bm{x}}{\|\Sigma^{-1}\bm{x}\|_{q}^{q-1}}\right].$
Then, we have that
$\sqrt{n}\left(\widehat{C}-\Sigma^{-1}\right)\Rightarrow-2\eta\Sigma^{-1}\mathbb{E}\left[\frac{\left(\mathrm{sign}(\Sigma^{-1}\bm{x})\odot|\Sigma^{-1}\bm{x}|^{q-1}\right)\bm{x}^{\top}}{\|\Sigma^{-1}\bm{x}\|_{q}^{q-1}}\right]\Sigma^{-1}.$
A.6 Proof of Theorem 5.2
Proof.
We define
$\Phi_{n}(C)=-\log\mathrm{det}C+\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}C\bm{x}\right]+2\delta\sum_{j=1}^{d}\sum_{k=1}^{d}\left(\frac{1}{n}\sum_{i=1}^{n}|[\bm{x}_{i}]_{k}|\right)|[C]_{kj}|+\delta^{2}\|C\|_{1,1}.$
For $C\succ 0$, the function $\Phi_{n}(C)$ is convex.
Case 1: $\gamma=1/2$.
$W_{n}(\bm{U})\overset{\vartriangle}{=}$ $n\left(\Phi_{n}(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})-\Phi(\Sigma^{-1})\right)$
$=$ $-n\log\det(\Sigma^{-1}+\frac{1}{\sqrt{n}}\bm{U})+n\log\det(\Sigma^{-1})+\sqrt{n}\mathbb{E}_{\mathbb{P}_{n}}\left[\bm{x}^{\top}\bm{U}\bm{x}\right]$
$+2\eta\sqrt{n}\sum_{j=1}^{d}\sum_{k=1}^{d}\left(\frac{1}{n}\sum_{i=1}^{n}|[\bm{x}_{i}]_{k}|\right)\left(|[\Sigma^{-1}]_{kj}+\frac{1}{\sqrt{n}}[\bm{U}]_{kj}|-|[\Sigma^{-1}]_{kj}|\right)$
$+o(1).$
Based on the proof of Theorem 5.1, we can have that
$\sqrt{n}(\widetilde{C}-\Sigma^{-1})\Rightarrow\arg\min_{\bm{U}}W(\bm{U}),$
where
$W(\bm{U})$ $=-\mathrm{tr}(\bm{U}\mathbf{G})+\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma)$
$+2\eta\sum_{j=1}^{d}\sum_{k=1}^{d}\mathbb{E}[|[\bm{x}]_{k}|]\left([\bm{U}]_{jk}\operatorname{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)+|[\bm{U}]_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
Case 2: $\gamma>1/2$.
$W(\bm{U})$ $=-\mathrm{tr}(\bm{U}\mathbf{G})+\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma).$
Case 3: $0<\gamma<1/2$.
$W(\bm{U})$ $=\frac{1}{2}\mathrm{tr}(\bm{U}\Sigma\bm{U}\Sigma)$
$+2\eta\sum_{j=1}^{d}\sum_{k=1}^{d}\mathbb{E}[|[\bm{x}]_{k}|]\left([\bm{U}]_{jk}\operatorname{sign}([\Sigma^{-1}]_{jk})\mathbb{I}([\Sigma^{-1}]_{jk}\not=0)+|[\bm{U}]_{jk}|\mathbb{I}([\Sigma^{-1}]_{jk}=0)\right).$
Since we have $\bm{x}\sim\mathcal{N}(\bm{0},\Sigma)$, then we have $\bm{x}]_{k}\sim\mathcal{N}(\bm{0},[\Sigma]_{kk})$. In this way,
$\mathbb{E}[|[\bm{x}]_{k}|]=\sqrt{\frac{2[\Sigma]_{kk}}{\pi}}.$